OpenStudy (anonymous):

derivate f(x)=ln(x2+3), x0=l f'(x)=

OpenStudy (welshfella):

use the chain rule

OpenStudy (welshfella):

does that help?

OpenStudy (anonymous):

but i don;t know to do that

OpenStudy (welshfella):

you haven't used the chain rule?

OpenStudy (anonymous):

no, i want to help my friend

OpenStudy (anonymous):

with your help^^

OpenStudy (welshfella):

ln(x^2 + 3) is a compound function - function within a function so we can write it as f(g(x) where g(x) = x^2 + 3 so derivative of f(g(x) = f'(g)x) * g'(x) = 1/ (x^2 + 3) * 2x = 2x/ (x^2 + 3) - that is f'(x)

OpenStudy (welshfella):

i used the standard derivatives D (ln u) = 1/u. and D (x^n) = nx^(n-1)

OpenStudy (solomonzelman):

Are you clear with evertthing? What other examples of chain rule, or what? (Please reply, don't just sit silent. We will help you)